OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 4 to 9. If the perimeter of the larger rectangle is 54 inches, what is the perimeter of the smaller rectangle?

OpenStudy (mathstudent55):

Use a proportion with these ratios. The ratio between the perimeters is 4:9 or 4/9. Since the larger one is 54, that corresponds to 9 (bec 9 is greater than 4), so the second ratio is x:54 or x/54 Now set the ratios equal to each other. That is a proportion. Then solve the proportion.

OpenStudy (anonymous):

can i just have the answer?

OpenStudy (mathstudent55):

4/9 = x/54 9x = 4 x 54 9x = 4 x 9 x 6 divide both sides by 9 x = 4 x 6 x = 24

OpenStudy (anonymous):

THank you